Crew 202 Sol Summary Report 03-Jan-2019
Summary Title: After a low-power night
Author’s name: Cesare Guariniello
Mission Status: The crew is continuing work on all research project. The morning was spent in power conservation mode, during which half of the crew performed an EVA and the other half stayed in the habitat to work on various projects. The power is now nominal.
Sol Activity Summary: After a good morning stretching and some light exercises, half of the crew went on EVA #4 while the rest worked on projects (and some origami) in the habitat. The afternoon was spent resting and working on research projects, including microbiology, geology, and astronomy. The crew is doing great, and every little difficulty or tiredness is overcome with laughter and a cohesion among crew members which is a joy to the eyes!
Look Ahead Plan: Tomorrow we will have a geology EVA in areas of sand dunes and continue our projects.
Anomalies in work: None
Weather: Sunny and cold with melting snow on the ground
Crew Physical Status: In perfect health
EVA: The Executive Officer, Crew Engineer, and Crew Journalist had an EVA to the region North of Kissing Camel Range, to collect clay and shale samples and identify sandy regions.
Reports to be filed: Sol Summary, Operations Report, Greenhab Report, EVA report, EVA request, Geology Report, Journalist Report
Support Requested: None
Commander, MDRS 202